OpenStudy (anonymous):

Peform a trigonometric substitution on the following integral. You do NOT need to do the new (d theta) integral. ((ln(x sqrt(49-x^2)) 5))/(7-xsqrt(49-x^2))

OpenStudy (lgbasallote):

trig sub without theta? i don't get it

OpenStudy (anonymous):

actually just ignore that part...but thats what it says on my question

OpenStudy (anonymous):

is that "5" an exponent?

OpenStudy (anonymous):

+5

Parth (parthkohli):

If we have \(a^2 - x^2\) inside the square root, then substitute \(x = a\sin\theta\).

Parth (parthkohli):

If we have something in that form, I mean...^

Parth (parthkohli):

I find everything confusing without \(\LaTeX\).

OpenStudy (anonymous):

??

OpenStudy (anonymous):

is it this? (god, please say no!): \(\huge \int \frac{ln(x\sqrt{49-x^2}+5)}{7-x\sqrt{49-x^2}}dx \)

OpenStudy (lgbasallote):

\[\sqrt{49 - x^2}\] so your triangle looks like  so 7\cos \theta = \sqrt{49 - x^2}\] and \[7 \sin \theta = x\]